OpenStudy (anonymous):

solve by finding the range and domain: x=(y+2)^2

OpenStudy (jdoe0001):

hmm have you covered parabolas yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@alexistheking777 @LynFran

OpenStudy (anonymous):

@carina98 @CBARREDO1 @chancemorris123 @dreamstoogood @awlas15

OpenStudy (cbarredo1):

range y=−2+x√,−2−√x domain x=y^2+4y+4

OpenStudy (anonymous):

Okay, how did you get this?